# Perturbation-based Learning for Recurrent Neural Networks

### v1

@@ 2.5 Decorrelation of neural inputs
Decorrelating neural input allows for more efficient neural representation by reducing the redundancy in neural activity, leading to improved efficiency in learning and faster learning rates. This phenomenon has found support in both biological studies (Wiechert et al.,, 2010; Cayco-Gajic et al.,, 2017) and artificial neural network research (Desjardins et al.,, 2015; Huang et al.,, 2018; Luo,, 2017; Ahmad et al.,, 2022). Additionally, in the context of global reinforcement methods such as WP or NP, where weight updates introduce substantial noise, the addition of decorrelation proves beneficial as it makes neural network systems less sensitive to noise (Tetzlaff et al.,, 2012).
Given these benefits, we investigate the integration of a decorrelation mechanism into our networks, following the approach in the original work. In contrast to Dalm et al., (2023), we apply this mechanism solely to the hidden units. This involves the transformation of correlated hidden layer input $x_{t-1}$ into decorrelated input $x^{*}_{t-1}$ via a linear transform $x^{*}_{t-1}=Dx_{t-1}$ with $D$ the decorrelation matrix. The update of recurrent units in the network is in this case defined as $x_{t}=f\left(Au_{t}+Rx^{*}_{t-1}\right)$ and the resulting neural connectivity is depicted in Figure 3.
@@
The updates of the decorrelation weights are performed using a particularly efficient learning rule proposed by Ahmad et al., (2022), which aims at reducing the cross-correlation between the neural outputs. The update is given by $D\leftarrow D-\epsilon\Delta D$ with learning rate $\epsilon$ and update
@@ 2.6 Experimental validation
We evaluate the effectiveness of the described methods in training RNNs using a series of experiments encompassing several objectives. Firstly, we evaluate the performance of the networks using three standard machine learning benchmarks: Mackey-Glass time series prediction, the copying memory task and a weather prediction task. These tasks are commonly employed in the literature to evaluate the performance of RNNs and other time-series prediction models. Secondly, we assess the scalability of the considered networks when incorporating an increased number of units. Lastly, we investigate in more detail the functioning of the decorrelation mechanism. Five different runs with random seeds are carried out for each experiment. The mean, maximum and minimum of these runs are then depicted.
Throughout these evaluations, we assess the perturbation-based learning methods NP, WP, and ANP, alongside the gradient-based BP learning method in conjunction with the Adam optimizer (Kingma and Ba,, 2014) for comparative analysis. Subsequently, we enhance all the methods by incorporating the decorrelation mechanism previously described into the hidden units of our networks. The resulting extended methods are named decorrelated node perturbation (DNP), decorrelated weight perturbation (DWP), decorrelated activity-based node perturbation (DANP) and decorrelated backpropagation (DBP). Detailed hyperparameters for each experiment can be found in Appendix B.
@@ 3.1 Mackey-Glass time series task
The Mackey-Glass time series task is a classic benchmark for assessing the ability of neural networks to capture and predict chaotic dynamical systems. The data is a sequence of one-dimensional observations generated using the Mackey-Glass delay differential equations (Mackey and Glass,, 1977), resulting in nonlinear, delayed, and chaotic time series. We reproduce the setup of Voelker et al., (2019), where the model is tasked with predicting 15 time steps into the future with a time constant of 17 steps and a length sequence of 5000 time steps.
@@
Figure 4 depicts the results for the Mackey-Glass experiment. Figure 4a, visualizes 500 time steps of a synthetically generated Mackey-Glass time series. Additionally, we depict the predictions made by a BP-trained model both before and after training. This example serves to provide a visual understanding of the dataset used in our study. Figure 4b shows the performance during training over the train and test set for the different methods. Figure 4c shows the final performance, computed as the mean performance over the last 50 epochs, facilitating a quantitative comparison between methods.
@@ 3.2 Copying memory task
The copying memory task is another well-established task for evaluating the memorization capabilities of recurrent models/units. In this task, the model has to memorize a sequence of bits and return the exact same sequence after a specified delay period. We build on the setup of Arjovsky et al., (2016), where the sequence comprises 8 distinct bit values, with 1 extra bit serving to mark the delay period and another extra bit indicating to the network when to reproduce the sequence. In our experiments, we use a length sequence of 100 and a delay period equal to 10.
@@ 3.3 Weather prediction task
In contrast to the other benchmarks used in this paper, this task does relies on real-world rather than synthetically generated data. The dataset used in this task contains climatological data spanning 1,600 U.S. locations from 2010 to 2013, obtained from https://www.ncei.noaa.gov/data/local-climatological-data/. We build on the setup of Zhou et al., (2021), where each data point consists of one single target value to be predicted 1, 24 and 48 hours in advance, and various input climate features. In our specific configuration, we exclude duplicated features with different units of measurement (retaining Fahrenheit-measured features) and select the ‘Dry bulb’ feature, a synonym for air temperature, as the target variable. The training data encompasses the initial 28 months, while the last 2 months are used for testing the model’s performance.
Figure 6 depicts the results for this task. In Figure 6a, we present a visualization of 2000 time steps of the target feature, ‘Dry bulb’, from the weather dataset. Additionally, we depict the predictions 48 hours ahead made by a BP-trained model both before and after training. This example serves to provide a visual understanding of the dataset used in our study. In Figure 6b we present the performance during training over the train and test set for the different methods for 48-hours ahead prediction. Additionally, these figures present the final performance, computed as the mean performance over the last 50 epochs, facilitating a quantitative comparison between methods. In Appendix C, we also include the results for 1-hour ahead and 24-hours ahead predictions.
@@ 3.4 Scaling performance
Here, we investigate the scalability of the described methods in RNNs with a single hidden and increasingly larger number of units. We accomplish this by analyzing the final performance on the 1-hour ahead weather prediction for networks with differing numbers of hidden units, trained using various methods. The number of hidden units ranges from 100 to 3000.
Figure 5 shows the final performance over the train and test sets of the networks with different configurations. Each configuration was repeated using five random seeds, and the results were computed as the mean of the successful executions within the set of five runs.
Figure 7: Scalability. Each data point represents the mean of five runs. Solid circles denote stability across all runs, with the mean calculated from the five stable runs. Empty circles indicate instability in some runs, and the mean is computed solely from the remaining stable executions. The absence of data points indicates that all the runs were unstable.
@@ 3.5 Decorrelation results
Here, we investigate the functionality of the decorrelation mechanism within the networks. Using the parameters that yield optimal task-dependent performance, we compare the degree of correlation of the neural outputs between a network incorporating the decorrelation mechanism and one that does not. To achieve this, we visualize the mean squared correlation across epochs during training, employing the 1-hour ahead weather prediction task.
Figure 8 illustrates the degree of correlation within the hidden units. This represents the decorrelation loss, which is calculated as the mean squared off-diagonal values of the lower triangular covariance matrix, computed with the recurrent inputs, $x^{*}_{t}$, to the hidden units at each timestep during input presentation, across epochs.
@@ 4 Discussion
Backpropagation through time is the default algorithm to train RNNs. Like backpropagation, it relies on the computation and propagation of gradient for the weight updates. However, the need for unrolling the RNN over time makes this algorithm computationally demanding and memory-intensive, especially with long input sequences. Furthermore, the non-local nature of its updates and the requirement to backtrack through time once the input sequence concludes can pose some challenges in implementing this training method.
In this paper, we introduced a viable alternative for training RNNs. Instead of using explicit gradients, our method stochastically approximates the gradients via perturbation-based learning. To this end, we extended the activity-based node perturbation approach (Dalm et al.,, 2023), which approximates stochastic gradient descent more accurately than other perturbation-based methods, to operate effectively in the time domain using RNNs. Inspired by the original work, we also incorporated a decorrelating mechanism into our networks to further improve convergence.
The results of our extensive validation show similar performance, convergence time and scalability of ANP and DANP when compared to BP. Remarkably, our approach exhibits superior performance, convergence time and scalability compared to standard perturbation-based methods, NP and WP. These promising findings suggest that perturbation-based learning holds the potential to rival BP’s performance while retaining inherent advantages over gradient-based learning. Notably, the computational simplicity of our method stands out, especially compared to BP, which requires a specific second phase for error computation across networks and a dedicated circuit for error propagation. This simplicity is particularly advantageous for RNNs, eliminating the need for time-unrolling that significantly increases computational load.
@@
Our results also support the relevance of perturbation-based learning for neuroscience. These learning approaches have been proposed as a model of learning in the brain (Miconi,, 2017; Fiete et al.,, 2007; Fiete and Seung,, 2006), as it leverages intrinsic brain noise for synaptic plasticity (Faisal et al.,, 2008), incorporating the noise as a feature rather than an obstacle. These approaches also contain global reinforcement signals that align with neuromodulatory signals in biological neural networks (Schultz,, 1998; Doya,, 2002; Marder,, 2012; Brzosko et al.,, 2019), and lack pathways for computing and delivering specific errors to each neuron. All these features yield a high level of biological plausibility to our approach, which is now substantiated by effective learning over practical tasks. Moreover, the requirement for decorrelated neural outputs in our approach is also a feature observed in biological neural networks, believed to be implemented through neural inhibition (Chini et al.,, 2022; Ecker et al.,, 2010).
The local nature of the required computations also facilitates the deployment of these methods on neuromorphic hardware. Embracing noise as a mechanism for learning can be highly suitable in settings where the computational substrate shows a high degree of noise (Gokmen,, 2021). This even holds when the noise cannot be measured since ANP still functions when we compare two noisy passes rather than a clean and a noisy pass (Dalm et al.,, 2023). Additionally, the gradient-free nature of the approach becomes valuable in settings where the computational graph contains non-differentiable components as in spiking recurrent neural networks; a type of network where effective training methods are still under exploration (Tavanaei et al.,, 2019; Wang et al.,, 2020; Neftci et al.,, 2019).
Finally, it is important to realize thatfurther validation on very large neural networks is necessary. This step is usually a common challenge in assessing the efficacy of biologically motivated learning algorithms (Bartunov et al.,, 2018). Nevertheless, current promising results open the door to gradient-free training of RNNs, offering exciting prospects for future research and applications in artificial intelligence, neuroscience and neuromorphic computing.
@@ Appendix B Hyperparameter settings
The different hyperparameter that our system contains are:
  • Gaussian noise variance $\sigma^{2}$
  • Learning rate $\eta$ for the forward and recurrent weights
  • Learning rate $\epsilon$ for the decorrelation weights
  • Batch size
  • Number of hidden units
  • Number of epochs
In all experiments, certain hyperparameters are fixed, while others are fine-tuned to optimize performance for each learning algorithm on a specific dataset. This ensures that we make a fair comparison between different learning algorithms. The number of epochs is consistently set to 500, as most algorithms demonstrate convergence at this point. The number of hidden units is set to 1000 in the experiments performed over the weather prediction task and the Mackey-Glass dataset. The number of hidden units is set to 500 in the copying memory task. In the scalability experiment, the number of hidden units serves as a variable parameter to evaluate its impact on final performance and stability. For experiments on the Mackey-Glass dataset, the batch size is 10, for the copying memory task, it is 1, and for the weather dataset, it is 50.
The Gaussian noise variance, $\sigma^{2}$, was consistently set to $10^{-2}$ in perturbation-based learning experiments, as the optimal value within the range [$10^{-1}$, $10^{-4}$]. Learning rates for forward/recurrent and decorrelation weights vary across experiments, chosen within the range [$10^{-1}$, $10^{-10}$]. Tables 1 and 2 present the specific learning rates employed for each dataset in conjunction with each learning algorithm. Learning rates exceeding the values mentioned above yielded unstable runs, characterized by weights becoming nan due to exploding values, or optimization getting stuck in local minima, resulting in inferior final performance. Conversely, smaller values result in slower learning, leading to inferior final performance. The backpropagation algorithm was employed in combination with the Adam optimizer using the default parameters.
Table 1: Learning rate $\eta$ for the forward and recurrent weights.
@@
Mackey-Glass $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $3\mathrm{e}{-6}$ $3\mathrm{e}{-5}$ $5\mathrm{e}{-7}$ $5\mathrm{e}{-7}$
Weather prediction $1\mathrm{e}{-2}$ $1\mathrm{e}{-2}$ $1\mathrm{e}{-3}$ $1\mathrm{e}{-3}$ $3\mathrm{e}{-5}$ $1.5\mathrm{e}{-4}$ $5\mathrm{e}{-7}$ $1\mathrm{e}{-6}$
Copying memory task $5\mathrm{e}{-2}$ $5\mathrm{e}{-2}$ $1\mathrm{e}{-3}$ $1\mathrm{e}{-3}$ $4\mathrm{e}{-5}$ $4\mathrm{e}{-5}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Table 2: Learning rate $\epsilon$ for the decorrelation weights.
NP DNP WP DWP ANP DANP BP DBP
Mackey-Glass - $5\mathrm{e}{-9}$ $5\mathrm{e}{-9}$ $5\mathrm{e}{-9}$ $5\mathrm{e}{-7}$
Weather prediction - $1\mathrm{e}{-7}$ $1\mathrm{e}{-7}$ $5\mathrm{e}{-7}$ $1\mathrm{e}{-7}$
Copying memory task - $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Appendix C Additional weather prediction experiments

### v3

@@ 2.5 Decorrelation of neural inputs
Decorrelating neural input allows for more efficient neural representation by reducing the redundancy in neural activity, leading to improved efficiency in learning and faster learning rates. This phenomenon has found support in both biological studies (Wiechert et al.,, 2010; Cayco-Gajic et al.,, 2017) and artificial neural network research (Desjardins et al.,, 2015; Huang et al.,, 2018; Luo,, 2017; Ahmad et al.,, 2022). It has also been shown that it significantly accelerates the training of deep neural networks (Dalm et al.,, 2024) by aligning SGD updates with the natural gradient through the use of uncorrelated input variables (Desjardins et al.,, 2015). Additionally, in the context of global reinforcement methods such as WP or NP, where weight updates introduce substantial noise, the addition of decorrelation proves beneficial as it makes neural networks less sensitive to noise (Tetzlaff et al.,, 2012).
Given these benefits, we investigate the integration of a decorrelation mechanism into our networks, presented in Dalm et al., (2023). We apply this mechanism solely to the hidden units, as the inputs fed into the network are low-dimensional, thus decorrelating them would not yield any noticeable changes and would increase computational costs. This involves the transformation of correlated hidden layer input $x_{t-1}$ into decorrelated input $x^{*}_{t-1}$ via a linear transform $x^{*}_{t-1}=Dx_{t-1}$ with $D$ the decorrelation matrix. The update of recurrent units in the network is in this case defined as $x_{t}=f\left(Au_{t}+Rx^{*}_{t-1}\right)$ and the resulting neural connectivity is depicted in Figure 3.
@@
The updates of the decorrelation weights are performed using a particularly efficient learning rule proposed by Ahmad et al., (2022), which aims at reducing the cross-correlation between the neural outputs. The update is given by $D\leftarrow D-\epsilon\Delta D$ with learning rate $\epsilon$ and update defined in Equation 6.
@@ 2.6 Experimental validation
We evaluate the effectiveness of the described methods in training RNNs using a series of experiments encompassing several objectives. Firstly, we evaluate the performance of the networks using three standard machine learning benchmarks: Mackey-Glass time series prediction, the copying memory task and a weather prediction task. These tasks are commonly employed in the literature to evaluate the performance of RNNs and other time-series prediction models. Secondly, we assess the scalability of the considered networks when incorporating an increased number of units. Lastly, we investigate in more detail the functioning of the decorrelation mechanism. Five different runs with random seeds are carried out for each experiment. The averages of these runs are then depicted with error bars indicating maximum and minimum values.
Throughout these evaluations, we assess the perturbation-based learning methods NP, WP, and ANP, alongside the gradient-based BP learning method in conjunction with the Adam optimizer (Kingma and Ba,, 2014) for comparative analysis. Please note that for simplicity, we use the term BP in this manuscript to refer to both backpropagation and its time-domain application, backpropagation through time (BPTT) (Werbos,, 1990), as they are essentially the same algorithm, with the distinction being the unrolling of the network over time. Similarly, the terms NP, WP, and ANP are used in this context to refer to their application in the time domain. Subsequently, we enhance all the methods by incorporating the decorrelation mechanism previously described into the hidden units of our networks. The resulting extended methods are named decorrelated node perturbation (DNP), decorrelated weight perturbation (DWP), decorrelated activity-based node perturbation (DANP) and decorrelated backpropagation (DBP). Detailed hyperparameters for each experiment can be found in Appendix B.
Our networks and experimental setups are developed using the programming language Python (Van Rossum et al.,, 1995) and the deep learning framework PyTorch (Paszke et al.,, 2019). The BPTT algorithm is implemented using the automatic differentiation tool from PyTorch. For reproducibility and to access the implementation code, please visit our GitHub repository 111 https://github.com/jesusgf96/NP_RNNs..
@@ 3.1 Mackey-Glass time series task
The Mackey-Glass time series task is a classic benchmark for assessing the ability of neural networks to capture and predict chaotic dynamical systems. The data is a sequence of one-dimensional observations generated using the Mackey-Glass delay differential equations (Mackey and Glass,, 1977), resulting in a nonlinear, delayed, and chaotic time series. We reproduce the setup of Voelker et al., (2019), where the model is tasked with predicting 15 time steps into the future with a time constant of 17 steps and a sequence length of 5000 time steps.
@@
Figure 4 depicts the results for the Mackey-Glass experiment. Figure 4a visualizes 500 time steps of a synthetically generated Mackey-Glass time series. Additionally, we depict the predictions made by a BP-trained model both before and after training. This example provides a visual understanding of the dataset used in our study. Figure 4b shows the performance during training over the train and test set for the different methods. Figure 4c shows the final performance, computed as the mean performance over the last 50 epochs, facilitating a quantitative comparison between methods.
@@ 3.2 Copying memory task
The copying memory task is another well-established task for evaluating the memorization capabilities of recurrent models/units. In this task, the model must memorize a sequence of bits and return the exact same sequence after a specified delay period. We build on the setup of Arjovsky et al., (2016), where the sequence comprises 8 distinct bit values, with 1 extra bit serving to mark the delay period and another extra bit indicating to the network when to reproduce the sequence. In our experiments, we use a length sequence of 100 and a delay period equal to 10.
@@ 3.3 Weather prediction task
In contrast to the other benchmarks used in this paper, this task relies on real-world rather than synthetically generated data. The dataset used in this task contains climatological data spanning 1,600 U.S. locations from 2010 to 2013.222Data can be obtained from https://www.ncei.noaa.gov/data/local-climatological-data/. We build on the setup of Zhou et al., (2021), where each data point consists of one single target value to be predicted 1, 24 and 48 hours in advance, and various input climate features. In our specific configuration, we exclude duplicated features with different units of measurement (retaining Fahrenheit-measured features) and select the ‘Dry bulb’ feature, which is a synonym for air temperature, as the target variable. The training data encompasses the initial 28 months, while the last 2 months are used for testing the model’s performance.
@@
Figure 6 depicts the results for this task. In Figure 6a, we present a visualization of 2000 time steps of the target feature, ‘Dry bulb’, from the weather dataset. Additionally, we depict the predictions 48 hours ahead made by a BP-trained model both before and after training. This example serves to provide a visual understanding of the dataset used in our study. In Figure 6b we present the performance during training over the train and test set for the different methods for 48-hours ahead prediction. Additionally, these figures present the final performance, computed as the mean performance over the last 50 epochs, facilitating a quantitative comparison between methods. We also include the results for 1-hour ahead and 24-hours ahead predictions in Appendix E.
@@ 3.4 Scaling performance
Here, we investigate the scalability of the described methods in RNNs with a single hidden layer and an increasingly larger number of units. We accomplish this by analyzing the final performance on the 1-hour ahead weather prediction task for networks with differing numbers of hidden units, trained using various methods. We chose this dataset as it is the most challenging among the considered datasets in this paper. The number of hidden units ranges from 100 to 3000.
Figure 7: Scalability. Panels show final train and test loss. Each data point represents the mean of five runs. Solid circles denote stability across all runs, with the mean calculated from the five stable runs. Empty circles indicate instability in some runs, and the mean is computed solely from the remaining stable executions. The absence of data points indicates that all the runs were unstable.
Figure 7 shows the final performance over the train and test sets of networks with different configurations. Each configuration was repeated using five random seeds, and the results were computed as the mean of the successful executions within the set of five runs.
@@ 3.5 Decorrelation results
Here, we investigate the functionality of the decorrelation mechanism within the networks. Using the parameters that yield optimal task-dependent performance, we compare the degree of correlation in the neural outputs between a network incorporating the decorrelation mechanism and one that does not. To achieve this, we visualize the mean squared correlation across epochs during training on the 1-hour ahead weather prediction task.
Figure 8 illustrates the degree of correlation within the hidden units. This represents the decorrelation loss, calculated as the mean squared off-diagonal values of the lower triangular covariance matrix, computed with the recurrent inputs, $x^{*}_{t}$, to the hidden units at each timestep during input presentation, across epochs.
@@ 4 Discussion
Backpropagation through time is the default algorithm to train RNNs. Like backpropagation, it relies on the computation and propagation of gradients for weight updates. However, the need to unroll the RNN over time makes this algorithm computationally demanding and memory-intensive, especially with long input sequences. Furthermore, the non-local nature of its updates and the requirement to backtrack through time once the input sequence concludes can pose some challenges in implementing this training method. In this paper, we introduced a viable alternative for training RNNs. Instead of using explicit gradients, our method approximates the gradients stochastically via perturbation-based learning. To this end, we extended the decorrelated activity-based node perturbation approach (Dalm et al.,, 2023), which approximates stochastic gradient descent more accurately than other perturbation-based methods, to operate effectively in the time domain using RNNs.
The results of our extensive validation show similar performance, convergence time and scalability of ANP and DANP when compared to BP. Remarkably, our approach exhibits superior performance, convergence time and scalability compared to standard perturbation-based methods, NP and WP. These promising findings suggest that perturbation-based learning holds the potential to rival BP’s performance while retaining inherent advantages over gradient-based learning. Notably, the computational simplicity of our method stands out, especially compared to BP, which requires a specific second phase for error computation across networks and a dedicated circuit for error propagation. This simplicity is particularly advantageous for RNNs, eliminating the need for time-unrolling that significantly increases computational load. Additional experiments, detailed in Appendix C, empirically validate our claims that perturbation-based methods are computationally simpler and less memory-intensive than the standard gradient-based backpropagation method. Note that other gradient-free learning methods for RNNs have been proposed that do not require separate learning passes, such as the random feedback local online (RFLO) algorithm (Murray,, 2019). We show that our method still compares favourably in terms of learning efficiency and computational cost in Appendix D. In follow-up work, further validation on very large neural networks is necessary. This step is usually a common challenge in assessing the efficacy of alternative learning algorithms (Bartunov et al.,, 2018). Other avenues for further research include exploring performance in more complex gated recurrent structures like LSTM and GRU units, as well as further validation on other challenging real-world datasets.
@@
Our approach has several limitations regarding biological fidelity. Our neural network models are significantly simplified, lacking spikes, structural plasticity, and detailed temporal dynamics and neural structure. However, our results are an important step in developing more biologically realistic learning approaches, while performing similarly to backpropagation, which is challenging to implement in biological circuits (Lillicrap et al.,, 2020; Whittington and Bogacz,, 2019). Numerous methods have been proposed for learning synaptic weights in artificial neural networks as alternatives, such as feedback alignment (Lillicrap et al.,, 2016; Nøkland,, 2016), target propagation (Bengio,, 2014; Lee et al.,, 2015; Ahmad et al.,, 2020; Bengio,, 2020), dendritic error propagation (Guerguiev et al.,, 2017; Sacramento et al.,, 2018), and spiking implementations combining global and local errors (Bellec et al.,, 2020). However, these approaches generally still require computing parameter or neuron-specific errors, necessitating specific and complex circuits for error propagation. In contrast, our approach uses global errors akin to neuromodulatory signals in biological neural networks (Schultz,, 1998; Doya,, 2002; Marder,, 2012; Brzosko et al.,, 2019). These global errors are uniformly spread across the network, eliminating the need for pathways to compute and deliver specific errors to each neuron. A key distinction of perturbation-based learning approaches, like ours, is their use of intrinsic brain noise for synaptic plasticity (Faisal et al.,, 2008), viewing noise as a beneficial feature rather than an obstacle. This mirrors the biological principle of using noise as a mechanism for learning, crucial for adapting to dynamic and unpredictable environments. Finally, our approach’s active decorrelation method is similar to decorrelation in the brain, thought to be implemented through neural inhibition (Ecker et al.,, 2010; Chini et al.,, 2022). This process reduces correlations among neurons, promotes efficient information coding, enhances learning efficiency, and mirrors the adaptive processes found in biological neural networks.
While not fully biologically realistic, our models have sufficient fidelity to be highly relevant for neuromorphic computation. The local nature of the required computations, combined with a global learning signal, facilitates the deployment of these methods on neuromorphic hardware (Sandamirskaya et al.,, 2022; Paredes-Vallés et al.,, 2024), devices that rely on distributed, localized processing units that mimic biological neurons. Embracing noise as a mechanism for learning can be highly suitable in settings where the computational substrate shows a high degree of noise (Gokmen,, 2021). This even holds when the noise cannot be measured since ANP still functions when we compare two noisy passes rather than a clean and a noisy pass (Dalm et al.,, 2023). This resilience to noise underscores the method’s suitability for real-world neuromorphic devices where noise is ubiquitous and difficult to control. Additionally, the gradient-free nature of the approach becomes valuable in settings where the computational graph contains non-differentiable components as in spiking recurrent neural networks; a type of network where effective training methods are still under exploration (Tavanaei et al.,, 2019; Wang et al.,, 2020; Neftci et al.,, 2019).
Concluding, our present findings open the door to efficient gradient-free training of RNNs, offering exciting prospects for future research and applications in artificial intelligence, neuroscience and neuromorphic computing.
@@ Appendix B Hyperparameter settings
In all experiments, certain hyperparameters remain fixed while others are fine-tuned to optimize the performance of each learning algorithm on specific datasets. This approach ensures a fair comparison between different algorithms. The number of epochs, hidden units, batch size, and Gaussian noise variance were kept consistent across learning algorithms for the various datasets, as detailed in Table 1. The value selected for the Gaussian noise variance is the optimal value within the range $[10^{-1},10^{-4}]$.
Table 1: Hyperparameters and their values used in the different experiments. Asterisks indicate that the value differs for each learning algorithm.
Hyperparameter Description Value per experiment
Mackey-Glass Copying Weather
$\sigma^{2}$ Gaussian noise variance $10^{-2}$ $10^{-2}$ $10^{-2}$
$\eta$ Learning rate for forward and recurrent weights * * *
$\epsilon$ Learning rate for decorrelation weights * * *
$N$ Number of hidden units 1000 500 1000
batch Batch size 10 1 10
epochs Number of training epochs 500 500 500
In contrast, the learning rates $\eta$ and $\epsilon$ varied for each algorithm depending on the dataset, as detailed in Table 2. The values for the learning rates are the optimal values within the range $[10^{-1},10^{-10}]$. Learning rates exceeding the depicted values yielded unstable runs, characterized by weights becoming undefined due to exploding values, or optimization getting stuck in local minima, resulting in inferior final performance. Conversely, smaller values resulted in slower learning, leading to inferior final performance. The backpropagation algorithm was employed in combination with the Adam optimizer using the default parameters.
Table 2: Learning rates $\eta$ and $\epsilon$ for the different learning algorithms across datasets. Dashes indicate that the parameter is not used.
@@
Mackey-Glass $\eta$ $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $5\mathrm{e}{-4}$ $3\mathrm{e}{-6}$ $3\mathrm{e}{-5}$ $5\mathrm{e}{-7}$ $5\mathrm{e}{-7}$
$\epsilon$ $5\mathrm{e}{-9}$ $5\mathrm{e}{-9}$ $5\mathrm{e}{-9}$ $5\mathrm{e}{-7}$
Copying $\eta$ $5\mathrm{e}{-2}$ $5\mathrm{e}{-2}$ $1\mathrm{e}{-3}$ $1\mathrm{e}{-3}$ $4\mathrm{e}{-5}$ $4\mathrm{e}{-5}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
$\epsilon$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Weather $\eta$ $1\mathrm{e}{-2}$ $1\mathrm{e}{-2}$ $1\mathrm{e}{-3}$ $1\mathrm{e}{-3}$ $3\mathrm{e}{-5}$ $1.5\mathrm{e}{-4}$ $5\mathrm{e}{-7}$ $1\mathrm{e}{-6}$
$\epsilon$ $1\mathrm{e}{-7}$ $1\mathrm{e}{-7}$ $5\mathrm{e}{-7}$ $1\mathrm{e}{-7}$
Appendix C Computational and memory costs analysis
Here, we profile the computational and memory costs of various learning algorithms considered in this paper. The analysis is conducted using the copying memory task, as it is the most studied and standard dataset in the literature. We measure the GPU operations time (in milliseconds) and allocated GPU memory (in bytes) for each algorithm on a single sample, presented one time, utilizing the PyTorch Profiler333 https://pytorch.org/docs/stable/profiler.html.. To account for potential hardware and system variability, such as GPU load, background processes, and memory management, each algorithm is analyzed five times, and the results are averaged. The GPU used for this analysis is the Nvidia Tesla T4444 https://images.nvidia.com/aem-dam/en-zz/Solutions/design-visualization/technologies/turing-architecture/NVIDIA-Turing-Architecture-Whitepaper.pdf.. The results of this analysis are presented in Figure 10.
Figure 10: Computational and memory costs. Each data point represents the average of five executions. The analysis is conducted on a single sample from the copying memory dataset, presented only once. The left plot uses a normal scale, while the right plot employs a logarithmic scale for better visualization.
Our empirical findings align with what is well-established in the literature: perturbation-based learning approaches are significantly less demanding in terms of computational and memory requirements compared to the standard gradient-based method, backpropagation. The GPU operations time is substantially lower for all perturbation-based methods considered in this study, with node perturbation being the most efficient, followed closely by weight perturbation and activity-based node perturbation. The allocated GPU memory remains comparable across perturbation-based methods, while it increases exponentially in backpropagation, particularly when incorporating the decorrelation mechanism. Unrolling the network over time with decorrelation steps results in substantial memory overhead, as all intermediate steps of all variables must be stored for each time step, unlike in perturbation-based methods.
Appendix D Comparison to another gradient-free method
To further extend the comparison of our approach with other gradient-free methods for training RNNs, we selected random feedback local online (RFLO). RFLO shares striking similarities with our approach and demonstrates highly efficient learning coupled with computational efficiency. Moreover, it employs eligibility traces and local updates in an online setup.
The setup for the RNN model used for the RFLO is the same as the one we employed, with the only difference being that the state of the hidden neurons evolves using a time constant $\tau$. This, while allowing for a more powerful temporal integration, adds computational complexity and leads to slower neural responses. The hidden units are defined as
$x_{t}=\left(\frac{1}{\tau}\right)f(Au_{t}+Rx_{t-1})+\left(1-\frac{1}{\tau}\right)x_{t-1}.$ (7)
The weight updates are defined as
$\Delta A$ $=\varepsilon_{t}x_{t},\,,\quad\Delta B=\left[B\varepsilon_{t}\right]p_{t},\,,\quad\Delta R=\left[B\varepsilon_{t}\right]q_{t},$ (8)
where $B$ denotes fixed random feedback weights, $\varepsilon_{t}$ denotes a local error defined as $y_{t}-y^{*}_{t}$, and $p_{t}$ and $q_{t}$ are eligibility traces capturing previous neural states defined as
$p_{t}$ $=\left(\frac{1}{\tau}\right)f^{\prime}(a_{t})x_{t-1}+\left(1-\frac{1}{\tau}\right)p_{t-1}$ (9)
$q_{t}$ $=\left(\frac{1}{\tau}\right)f^{\prime}(a_{t})u_{t-1}+\left(1-\frac{1}{\tau}\right)q_{t-1}$ (10)
In our experiments, the parameter $\tau$ is set to 10 as in the original work, and the learning rate is empirically selected to be the optimal value in the task within the range $[10^{-1},\ldots,10^{-10}]$, with $10^{-5}$ being the optimal value.
For simplicity, we focused on the most effective perturbation-based approaches, ANP and DANP, and the gradient-based baselines, BP and DBP. The learning efficiency comparison was conducted on the 48-hour weather prediction task, which is the most challenging task in our study. The results of this comparison are presented in Figure 11.
a)
b)
Figure 11: 48-hour ahead weather prediction results including RFLO. a) Performance during training over the train and test set for the different methods, represented in a logarithmic scale. b) Performance for the different methods, computed as the mean performance over the last 50 epochs.
We observe that RFLO learns the task efficiently but lags slightly behind other methods in terms of convergence speed. While its final training performance is comparable to the other approaches, its generalization capability (in terms of the test loss) is much lower.
The computational cost analysis is performed following the procedure described in Section C of this document. The results of this analysis are shown in Figure 12.
Figure 12: Computational and memory costs. Each data point represents the average of five executions. The analysis is conducted on a single sample from the copying memory dataset, presented only once. The left plot uses a normal scale, while the right plot employs a logarithmic scale for better visualization.
RFLO is slightly more efficient in terms of operation time but slightly less efficient in terms of memory usage.
Appendix E Additional weather prediction experiments